### v1

@@ 3.1.3 Cleaning the Data
After processing the entities and themes, the next step is to clean redundant data. We consider two types of redundant articles: duplicated or remarkably similar and very short. There are various reasons for article duplication. For example, the same article from the same source could be parsed again in the given dataset, or the same article was obtained from a different source. We decided to remove articles with the same themes, entities, and tones (provided with GCAM), as articles we have already processed to mitigate over-representation in data of the same popular articles.
To handle the redundant short article, we removed all the articles containing less than a given predefined threshold of valid entities and themes.
@@ 3.1.4 Splitting the Data
Lastly, we deployed a splitting mechanism after completing the previous processing tasks. Given a time frame, it will split the given dataset of articles into smaller datasets, such as daily and monthly separated datasets. We perform this for several reasons: First, the embedding representation of entities and articles should based on the relevant reality - for example, currently, we expect Joe Biden’s embedding representation to be closer to the US President entity than Donald Trump’s embedding. Moreover, to keep the method lightweight, we consider only part of the dataset at all times if necessary. During this study, we noticed that most events have a short life period. Therefore, the information loss can be contained with the predefined time-frame split.
@@ 3.2 Embedding Generation for Entities
This stage of the method is responsible for producing embedding representations for entities. Given time-separated datasets of processed entities found in articles, this stage generates embedding for each entity. To perform this, we train GloVe models for each time-partitioned dataset, based on the entities co-occurrences in the articles. To improve the generation of time-sensitive embeddings for entities, we developed a pooling mechanism of additional articles before the training process to represent current and historical information.
@@ 3.2.1 Embedding Generation for Entities - Adapting GloVe
$X_{ij}$ is the number of times entity $i$ occurs in the same context window (article) as entity $j$;
$X_{i}=\sum_{j}X_{ij}$;
$P_{ij}=\frac{X_{ij}}{X_{i}}$; is the probability that entity $j$ appears in the same article as the entity $i$.
@@
Before training the GloVe model, we developed a pooling mechanism to ensure that the entities’ embeddings generated in a specific period contain recent and historical information. The pooling mechanism was constructed as follows: First, we sampled articles from previous periods in a decreasing manner, adding them to the current period corpus used to train the model. It is important to note that we did not sample data from the future to prevent leakage or bias to the entities’ embeddings. To define the pooling mechanism we use the following notations: The dataset of all the processed articles’ entities that appeared in a set time frame $t$ is marked as $M_{t}$. $rand(M_{k},p)$ is a function that randomly selects articles’ entities without replacement with $p$ probability to pick an article. The article corpus for each time frame was assembled in the following manner:
$Corpus_{M_{k}}=\{{M_{k}},rand(M_{i-1},p_{1}),...,rand(M_{n},p_{n-1})\}$
@@ 3.3 Embedding Generation for Articles
We utilized two distinct techniques to create news article embeddings. Firstly, we deploy a sentence embedding model on the embedded entities space so that given a set of embedding representations of article entities, this model generates a vector representing the article. Secondly, we train a semi-supervised DNN, similar to Ohi et al. work [60]; we did so on the tagged article event data. This model receives input from the naive articles embedding to produce a new vector space that considers the cross-document relationships given shared events. After completing the two approaches, we generate a third and final representation for the article embeddings - where we rely on the synergy of the two previous methods.
In the following sections, we describe the steps required to generate our method’s final article embedding: deploy sentence embedding model (see Section 3.3.1), then train a DNN semi-supervised model (see Section 3.3.2). Lastly, we perform a synergy based on the previous steps (see Section 3.3.3).
@@ 3.3.1 Embedding Generation for Articles Using SIF
In this stage, we construct a naive article embedding representations using the SIF algorithm [27]. Conceptually, to represent a complex object one can apply a simple aggregator function on the object’s component embeddings. By this logic, the SIF model receives as input a list of the entities’ embeddings that appeared in each article, as generated in the previous step (see Section 3.2), and outputs article embeddings.
We chose the SIF model as the aggregator function [27] due to its simplicity and competitive performance. The underlying intuition is that the embeddings of too frequent words should be down-weighted when summed with those of less frequent ones is compatible with the use-case of entities in articles - frequent entities provide less context required for distinguishing articles from one another.
We implemented this and considered it as the base approach for the task of article embeddings. However, there might be a drawback to such an approach, it might suffer from underfitting. The simplicity of the SIF model does not account for additional information given by a group of certain entities. To uncover that pattern one can use the information between articles - in our case articles that discuss common events.
@@ 3.3.2 Embedding Generation for Articles Using DNN Semi Supervised
We chose the Siamese Network model for this task, introduced by Bromley et al. [61], based on two identical Deep Neural Networks. The central concept of the approach is independent of the complexity of the model. Therefore, we chose a basic architecture for the models, which is independent of the specific time frame of the training data. This means that we applied the same model architecture to the entire dataset. The input layer of the model equals the size of the embedding generated by the SIF model. Three hidden layers are densely connected. The output layer of the model equals the input embedding size. The ReLU activation functions were configured on top of each layer, with uniform kernel initialize functions and L2 regularization functions. Figure 3 presents an overview of the suggested implementation of the Siamese Network.
@@
Our aim in using such network architecture is to generate article embeddings. For our problem, we define the anchor as an article and we consider an article that shares common events for the anchor as positive and an article that does not consider a negative sample respectively.
@@
Many works have discussed the benefit of hard negative mining in constructing triplets that produce useful gradients and therefore help triplet loss networks converge quickly [63, 64]. However, previous works also show that hard positive examples can increase clustering within a class [65]. Therefore, policy within each batch and optimized our loss function using ADAM optimizer [66] - we first sample $N$ triplet pairs, and out of those we select the top $K_{1}$ most violated matches in each sample. This is done by computing the pairwise distances between all articles within the initial batch. The pairwise distances were calculated as follows:
@@ 3.3.4 Articles Embedding Evaluation
As explained in the previous sections, each method generated an embedding representation for each article by receiving the article’s entities embedding as input. We perform extensive experiments to evaluate the performance of the proposed methods for event-centric article embedding. We measure the methods’ performances on the following task: Given a periodic dataset of articles’ embeddings and their tagged events, we create pairwise permutations of all the articles in the given dataset. A pair of articles is considered close or instead tagged as positive if both articles share at least a single joint event. A simple Euclidean distance between each pair’s embedding representation concludes their ranking score.
To validate the performance of our method, we performed a cross-dataset evaluation on different periods across the collected data without fine-tuning the models with the variant training data. The evaluation process considered the three proposed methods for article embeddings. To measure the methods we consider the task of common event attribution, and evaluate two metrics: precision-recall AUC, and ROC AUC [68]. Moreover, we also conducted the Friedman test [69] and Nemenyi post-hoc [70] to determine whether a significant statistical difference exists between the article embedding generation methods.
@@ 4 Experimental Setup
We evaluated our methods (see Section 3 and Figure 1) on monitored news sources by GDELT from 02/2015 to 01/2021. Finally, after the embedding generation for articles is complete, cosine similarity between all the article pairs in the same time frame is computed, to determine which articles share at least a single common event. The dataset is described in detail in Section 4.1, the experiment is described in Section 4.2, and the results are presented in Section 5.
@@ 4.1 Data Description
Throughout this paper, three different types of information are employed, including global knowledge graphs, events, and article-event mentions. These datasets are accessible by GDELT as three different types of files, for additional information on the GDLET source data (see A.2). The period of the data that have been collected is 02/2015 to 01/2021. Overall, in this research, we have evaluated 850,492 unique articles, 1,002,967 events, and 3,551,906 mentions.
We decided to collect all the articles from various popular media sources; each is viewed significantly more by people who identify as liberals or conservatives. We depended on previous works that classified different sources to a specific political view, such as described in a study by Iyengar et al. [71]. We have also decided to monitor the affiliates of those networks if they exist. For example, Fox News has more than 50 local affiliates, so we collected data on the articles from all of Fox News affiliates and classified them as the same source ’FoxFamily’. Including affiliates, we have gathered articles from 605 media sources (see A.2.1). Figure 4 presents the distribution of media sources in the collected data sets, labeled by political affiliation.
Figure 4: This figure maps each monitored media source to the articles published by it in the preprocessed dataset. Conservative and Liberal media sources are labeled in blue and red respectively.
@@
4.1.1 Labled Datasets Construction
We validate the proposed methods by performing cross-dataset testing by using our method. We have separated the entire collected article data into 66 datasets separated by months. We derived from the GDELT open-sourced data, Table LABEL:tab:datasets_statistics presents the characteristics of each monthly dataset included in the experiment. While Table 1 shows yearly aggregated statistical information of those datasets. It is important to note that some datasets were relatively smaller in the number of articles, events, and mentions of appearances. The difference in the datasets’ sizes led us to assume there was an inconsistency in the raw GDELT data.
@@ 4.1.1 Labled Datasets Construction
  • Monthly Mentioned Events: A large-scale dataset consists of every possible permutation of 2 articles in the same month. Given the embedding representation of each pair of articles, we calculate a prediction probability score that those articles shared at least a single joint event. It is worth noting that working with this dataset is challenging because it is highly imbalanced, and the probability of pairing articles with relatively far creation dates is substantially lower since events’ mentions appearances over time were skewed toward their creation date.
@@ 4.2 Experiments
In this section, we provide the implementation details and discuss the evaluation process of each step of the proposed methodology. For additional implementation notes, regarding the research experiments, see Section A.1.
@@
In the processing of the entities task, we define the following parameters:
@@
  • For determining whether an entity already appeared, the hamming distance similarity threshold was set to two. We have picked the threshold by manually selecting a few entities and searching their closest entities by hamming distance. We noticed that for the majority of cases, a hamming distance of two or less was caused by a typographical error.
  • To handle the redundant short articles, we removed all the articles containing less than four valid entities and themes. We assume that an article with such a few number of entities is irrelevant or there was an error in its entities extraction process.
Figure 5: Siamese-Network: Train Triplet Loss. The X-axis is the monitored training steps, every 4 steps the average loss was calculated. The labels represent each Siamese model that trained for a given month.
@@
The GloVe parameters were set as learning rate = 0.05, max loss = 10, and vector size = 175. We fixed these parameters for all the experiments.
We manually evaluated the generated GloVe embedding by doing the following: We selected 50 different themes and entities, overall. For each entity or theme, we searched for the five closest neighbors on a monthly trained GloVe model. We had two criteria for selecting each entity:
@@
After manually examining the selected entities and themes, we considered their relationships with others to be reasonable. For examples of the effect the pooling mechanism had on the entities embeddings (see Section A.3).
Embedding Generation for Articles. In each experiment, we split the events in the given dataset by a 70%-30% split to train and test by their order of appearance, as suggested by Nguyen et al. [72]. Nguyen et al. stated that ”The 70–30 split, widely employed in machine learning, strikes a balance between an adequate training set size and a sufficiently large test set, ensuring a robust evaluation of the model’s performance” [72]. In addition, data leakage articles duplication between the train and test set were removed.
We conducted this on each monthly dataset and evaluated suggested methods and the baseline. When training the Siamese Networks we defined fixed values for the hyper-parameters: ADAM optimizer with a Learning Rate of 0.0015. The Siamese Triplet model contained three DNN network architectures simple input and output layers of 175 units with three hidden layers of 128 units, we have also defined L2 regularization on top of each layer of the model. Each batch the size of 64, which consisted of 50 random pairings of articles and 14 hard pairings, as described by our defined triplets mining policy. [^11] The models were not trained on epochs but on 10,000 batches because of the sheer volume of the pair article’s permutations. We chose 10,000 batches since we saw a convergence across the datasets beforehand, which made sense in a relatively small network as we constructed. This can be exemplified by the train loss of the Siamese models we train in 2020 monthly datasets (see Figure 5). Each label is the relevant Siamese model in a given month. We constructed the embeddings, as defined in 3.3.2.
Performance Evaluation Figure 5 presents the training triplet loss curve of the Siamese network to visualize the training process on each dataset in 2020. We observed that the training loss of the models converges to a similar loss value.
@@ 5.1 Periodical Theme and Entity Embeddings
In this section, we present the number and type of each periodic embedding we generated in this work. As part of our study’s methodology, we have developed a new way of generating embeddings that are relevant to different periods. Specifically, we have created monthly embeddings for every valid entity or theme that we have identified as defined in Section 3.1. This means that even if a valid entity appears only once in all of the monitored media sources, in a given month, we can still create a monthly relevant embedding representation for it by using the GloVe model training process, as we described in Section 3.2. Our methodology involved monitoring various features from the news articles, including Person, Location, Date, Organization, and Theme. We constructed monthly embeddings for distinct entities from 02/2015-01/2021, and the count of these entities is presented in Table 2. Throughout the 66 months of monitoring, we aimed to generate 331,630 embeddings for news entities and themes each month.
Table 2: This table presents the number of unique generated embeddings per monitored entity type.
@@
Table 3: This Table contains the mean and standard deviation, of each method on both daily and monthly common event attribution tasks, for PR and ROC AUC metrics.
@@
max width= Method Monthly Evaluation Daily Evaluation Precision-Recall AUC ROC AUC Precision-Recall AUC ROC AUC SIF $0.348\pm 0.048$ $0.958\pm 0.015$ $0.547\pm 0.055$ $0.955\pm 0.014$ Siamese Network $0.344\pm 0.054$ $\mathbf{0.974\pm 0.012}$ $0.548\pm 0.060$ $\mathbf{0.972\pm 0012}$ Concatenation $\mathbf{0.369\pm 0.048}$ $\mathbf{0.974\pm 0.010}$ $\mathbf{0.572\pm 0.055}$ $\mathbf{0.972\pm 0.010}$

### v2

@@ 3.1.3 Cleaning the Data
After processing the entities and themes, the next step is to clean up redundant data. We identify two types of redundant articles: duplicates and very short articles. Article duplication can occur for several reasons, such as the same article being parsed multiple times from the same source or being obtained from different sources. To prevent over-representation of popular articles, we remove any articles that share the same themes, entities, and tones (provided by GCAM) as those already processed. To handle the redundant short article, we removed all the articles containing less than a given predefined threshold of valid entities and themes.
@@ 3.1.4 Splitting the Data
Lastly, after completing the previous processing tasks, we implemented a splitting mechanism. This mechanism divides the given dataset of articles into smaller subsets based on a specified time frame. We do this for several reasons: first, the embedding representation of entities and articles should reflect the relevant contextual news.
For instance, we currently expect Joe Biden’s embedding representation to be closer to the US President entity than Donald Trump’s embedding. Additionally, to keep the method lightweight, we consider only part of the dataset at a time if necessary. Our study showed that most events have a short lifespan, so any information loss can be minimized by using a predefined time-frame split.
@@ 3.2 Embedding Generation for Entities
This stage is responsible for producing embedding representations for entities. Given time-separated datasets of processed entities found in articles, this stage generates embedding for each entity. To perform this, we train GloVe models for each time-partitioned dataset based on the entity co-occurrences in the articles. To improve the generation of time-sensitive embeddings for entities, we developed a pooling mechanism for additional articles before the training process to represent current and historical information.
@@ 3.2.1 Embedding Generation for Entities - Adapting GloVe
$X_{ij}$ is the number of times entity $i$ occurs in the same context window (article) as entity $j$. We define $X_{i}$ as: $X_{i}:=\sum_{j}X_{ij}$, and $P_{ij}:=\frac{X_{ij}}{X_{i}}$ is the probability that entity $j$ appears in the same article as the entity $i$.
@@
Before training the GloVe model, we developed a pooling mechanism to ensure that the entities’ embeddings generated for a specific period contain both recent and historical information. The pooling mechanism was constructed as follows: first, we sampled articles from previous periods in decreasing amounts, adding them to the current period corpus used to train the model. It is important to note that we did not sample data from the future to prevent leakage or bias in the entities’ embeddings. The pooling mechanism is defined using the following notations: the dataset of all the processed articles’ entities that appeared in a set time frame $t$ is marked as $M_{t}$. We define $rand(M_{k},p)$ to be a function that randomly selects articles’ entities without replacement with $p$ probability to pick an article. The article corpus for each time frame was assembled in the following manner:
$Corpus_{M_{k}}:=\{{M_{k}},rand(M_{i-1},p_{1}),...,rand(M_{n},p_{n-1}).\}$
@@ 3.3 Embedding Generation for Articles
We utilized two distinct techniques to create news article embeddings. Firstly, we deploy a sentence embedding model on the embedded entities space so that, given a set of embedding representations of article entities, this model generates a vector representing the article. Secondly, we trained a semi-supervised DNN, similar to Ohi et al. work [59]; we did so on the tagged article event data. This model receives input from the naive article embedding to produce a new vector space that considers the cross-document relationships given shared events. After completing the two approaches, we generate a third and final representation for the article embeddings - where we rely on the synergy of the two previous methods.
In the following sections, we describe the steps required to generate our method’s final article embedding: deploy a sentence embedding model (see Section 3.3.1), then train a DNN semi-supervised model (see Section 3.3.2). Lastly, we perform a synergy based on the previous steps (see Section 3.3.3).
@@ 3.3.1 Embedding Generation for Articles Using SIF
In this stage, we construct a naive article embedding representations using the SIF algorithm [27]. Conceptually, to represent a complex object, one can apply a simple aggregator function on the object’s component embeddings. By this logic, the SIF model receives as input a list of the entities’ embeddings that appeared in each article, as generated in the previous step (see Section 3.2), and outputs article embeddings.
We chose the SIF model as the aggregator function [27] due to its simplicity and competitive performance. The underlying intuition is that the embeddings of too frequent words should be down-weighted when summed with those of less frequent ones, which is compatible with the use-case of entities in articles - frequent entities provide less context required for distinguishing articles from one another.
We implemented this and considered it as the base approach for the task of article embeddings. However, there might be a drawback to such an approach, it might suffer from underfitting. The simplicity of the SIF model does not account for additional information given by a group of certain entities. To uncover that pattern, one can use the information between articles - in our case articles that discuss common events.
@@ 3.3.2 Embedding Generation for Articles Using DNN Semi Supervised
We chose the Siamese Network model for this task, introduced by Bromley et al. [60], based on two identical DNNs. The central concept of the approach is independent of the complexity of the model. Therefore, we chose a basic architecture for the models, which is independent of the specific time frame of the training data. This means that we applied the same model architecture to the entire dataset. The input layer of the model equals the size of the embedding generated by the SIF model. Three hidden layers are densely connected. The output layer of the model equals the input embedding size. The ReLU activation functions were configured on top of each layer, with uniform kernel initialize functions and L2 regularization functions. Figure 3 presents an overview of the suggested implementation of the Siamese Network.
@@
Our goal in using this network architecture is to generate article embeddings. For our purposes, we define the anchor as an article. An article that shares common events with the anchor is considered a positive sample, while an article that does not is considered a negative sample.
@@
Many works have discussed the benefit of hard negative mining in constructing triplets that produce useful gradients and therefore help triplet loss networks converge quickly [62, 63]. However, previous works also show that hard positive examples can increase clustering within a class [64]. Therefore, we defined the following policy using the ADAM optimizer [65]: we first sample $N$ triplet pairs, and out of those, we select the top $K_{1}$ most violated matches in each sample. This is done by computing the pairwise distances between all articles within the initial batch. The pairwise distances were calculated as follows:
@@ 3.3.4 Articles Embedding Evaluation
As explained in the previous sections, each method generated an embedding representation for each article by receiving the article’s entity embedding as input. We perform extensive experiments to evaluate the performance of the proposed methods for event-centric article embedding. We measure the methods’ performances on the following task: Given a periodic dataset of articles’ embeddings and their tagged events, we create pairwise permutations of all the articles in the given dataset. A pair of articles is considered close or instead tagged as positive if both articles share at least a single joint event. A simple Euclidean distance between each pair’s embedding representation concludes their ranking score.
To validate the performance of our method, we performed a cross-dataset evaluation on different periods across the collected data without fine-tuning the models with the variant training data. The evaluation process considered the three proposed methods for article embeddings. To measure the methods, we consider the task of common event attribution and evaluate two metrics: precision-recall AUC and ROC AUC [67]. Moreover, we also conducted the Friedman test [68] and Nemenyi post-hoc [69] to determine whether a significant statistical difference exists between the article embedding generation methods.
@@ 4 Experimental Setup
We evaluated our methods (see Section 3 and Figure 1) on monitored news sources by GDELT from 02/2015 to 01/2021. Finally, after the embedding generation of articles is complete, cosine similarity between all the article pairs in the same time frame is computed to determine which articles share at least a single common event. The dataset is described in detail in Section 4.1, the experiments are described in Section 4.2, and the results are presented in Section 5.
@@ 4.1 Data Description
Throughout this paper, three different types of information are employed, including global knowledge graphs, events, and article-event mentions. These datasets are accessible by GDELT as three different types of files, for additional information on the GDLET source data (see Appendix A.2). The period of the data that have been collected is 02/2015 to 01/2021. Overall, in this research, we have evaluated 850,492 unique articles, 1,002,967 events, and 3,551,906 mentions.
We decided to collect all the articles from various popular media sources; each is viewed significantly more by people who identify as liberals or conservatives. We depended on previous works that classified different sources to a specific political view, such as described in a study by Iyengar et al. [70]. We have also decided to monitor the affiliates of those networks if they exist. For example, Fox News has more than 50 local affiliates, so we collected data on the articles from all of Fox News affiliates and classified them as the same source ’FoxFamily’. Including affiliates, we have gathered articles from 605 media sources (see Appendix A.2.1). Figure 4 presents the distribution of media sources in the collected data sets, labeled by political affiliation.
Figure 4: This figure maps each monitored media source to the articles published by it in the preprocessed dataset. Conservative and Liberal media sources are labeled in blue and red, respectively.
@@
4.1.1 Labeled Datasets Construction
We validate the proposed methods by performing cross-dataset testing by using our method. We have separated the entire collected article data into 66 datasets, separated by months. We derived from the GDELT open-sourced data, and Table LABEL:tab:datasets_statistics presents the characteristics of each monthly dataset included in the experiment. Table 1 shows yearly aggregated statistical information of those datasets. It is important to note that there is a variance in the datasets’ size. The difference in the datasets’ sizes led us to assume there was an inconsistency in the raw GDELT data.
@@ 4.1.1 Labeled Datasets Construction
  • Monthly Mentioned Events: A large-scale dataset consists of every possible permutation of 2 articles in the same month. Given the embedding representation of each pair of articles, we calculate a prediction probability score that those articles shared at least a single joint event. It is worth noting that working with this dataset is challenging because it is highly imbalanced, and the probability of pairing articles with relatively distant creation dates is substantially lower since events’ mentions appearances over time were skewed toward their creation date.
@@ 4.2 Experiments
In this section, we will outline the implementation details and discuss the evaluation process for each step of our proposed methodology. We conducted experiments to determine whether our approach improves the performance of common event attribution in news articles, whether this improvement is consistent across different data distributions, and whether our approach is independent of the initial embedding method. For additional implementation notes regarding the research experiments, see Appendix A.1.
@@
In the processing, we defined the following parameters:
@@
  • For determining whether an entity already appeared, the hamming distance similarity threshold was set to two. We have picked the threshold by manually selecting a few entities and searching for their closest entities by hamming distance. We noticed that, in the majority of cases, a hamming distance of two or less was caused by a typographical error.
  • To handle the redundant short articles, we removed all the articles containing less than four valid entities and themes. We assume that such article is irrelevant or that there was an error in its entity extraction process.
Figure 5: Siamese-Network: Train Triplet Loss. The X-axis is the monitored training steps; for every 4 steps, the average loss was calculated. The labels represent each Siamese model that was trained for a given month.
@@
The GloVe parameters were set as The GloVe parameters were set as learning rate = 0.05, max loss = 10, and vector size = 175. We fixed these parameters for all the experiments.
We manually evaluated the generated GloVe embedding by doing the following: We selected 50 different themes and entities overall. For each entity or theme, we searched for the five closest neighbors on a monthly trained GloVe model. We had two criteria for selecting each entity:
@@
After manually examining the selected entities and themes, we found their relationships with others to be reasonable. For examples of the pooling mechanism’s impact on the entities’ embeddings, see Appendix A.3.
Embedding Generation for Articles. In each experiment, we split the events in the given dataset by a 70%-30% ssplit to train and test by their order of appearance, as stated by Nguyen et al. ”The 70–30 split, widely employed in machine learning, strikes a balance between an adequate training set size and a sufficiently large test set, ensuring a robust evaluation of the model’s performance” [71]. In addition, data leakage articles duplication between the train and test set were removed.
When training the Siamese networks, we defined fixed values for the hyperparameters: ADAM optimizer with a learning rate of 0.0015. The Siamese Triplet model contained three DNN network architectures: simple input and output layers of 175 units, with three hidden layers of 128 units. We have also defined L2 regularization on top of each layer of the model. Each batch had a size of 64, which consisted of 50 random pairings of articles and 14 hard pairings, as described by our defined triplet mining policy. [^11] The models were not trained on epochs but on 10,000 batches because of the sheer volume of the pair article’s permutations. We chose 10,000 batches since we saw convergence across the datasets beforehand, which made sense in the relatively small network we constructed. This can be exemplified by the train loss of the Siamese models we train in 2020 monthly datasets (see Figure 5). Each label is the relevant Siamese model for a given month. We constructed the embeddings, as defined in 3.3.2.
Performance Evaluation To evaluate the performance of our approach to generating news embedding, we performed the following steps:
  1. 1.
    We deployed the semi-supervised event approach, SIF, and their concatenation – for all tasks of common event attribution.
  2. 2.
    We evaluated their performance on those tasks on all the collected datasets from GDELT, to determine whether the improvement was generalized over different data distributions.
  3. 3.
    We implemented a spaCy language model to generate embedding for articles given their themes and entities and evaluated them against our method.
  4. 4.
    Lastly, we deployed our semi-supervised approach on the spaCy models’ embeddings to verify that the approach is agnostic to the initial embedding method. This Siamese network was trained in the same manner as we described above.
@@ 5.1 Periodical Theme and Entity Embeddings
In this section, we present the number and type of each periodic embedding we generated in this work. As part of our study’s methodology, we have developed a new way of generating embeddings that are relevant to different periods. Specifically, we have created monthly embeddings for every valid entity or theme that we have identified, as defined in Section 3.1. This means that even if a valid entity appears only once in all the monitored media sources in a given month, we can still create a relevant monthly embedding representation for it by using the GloVe model training process, as we described in Section 3.2. Our methodology involved monitoring various features of the news articles, including Person, Location, Date, Organization, and Theme. We constructed monthly embeddings for distinct entities from 02/2015 to 01/2021, and the count of these entities is presented in Table 2. Throughout the 66 months of monitoring, we aimed to generate 331,630 embeddings for news entities and themes each month.
Table 2: The number of unique generated embeddings per monitored entity type.
@@
Table 3: The mean and standard deviation, of each method on both daily and monthly common event attribution tasks, for PR and ROC AUC metrics.
@@
max width= Method Monthly Evaluation Daily Evaluation Precision-Recall AUC ROC AUC Precision-Recall AUC ROC AUC SIF $0.348\pm 0.048$ $0.958\pm 0.015$ $0.547\pm 0.055$ $0.955\pm 0.014$ Siamese Network $0.344\pm 0.054$ $\mathbf{0.974\pm 0.012}$ $0.548\pm 0.060$ $\mathbf{0.972\pm 0012}$ Concatenation $\mathbf{0.369\pm 0.048}$ $\mathbf{0.974\pm 0.010}$ $\mathbf{0.572\pm 0.055}$ $\mathbf{0.972\pm 0.010}$ spaCy CPU Language Model $0.009\pm 0.003$ $0.877\pm 0.025$ $0.068\pm 0.020$ $0.876\pm 0.025$ Siamese Network - spaCy CPU Language Model $0.012\pm 0.004$ $0.933\pm 0.018$ $0.089\pm 0.028$ $0.932\pm 0.018$
@@
(a) Monthly Precision-Recall AUC Comparison
(b) Monthly ROC AUC Comparison
(c) Daily Precision-Recall AUC Comparison
(d) Daily ROC AUC Comparison
Figure 7: Statistical Analysis of Article Generation Using the spaCy Model - Comparison of performance using the Friedman and Nemenyi tests with a 5% significance level in the daily and monthly event attribution task.